OpenStudy (anonymous):

solve the equation: log2 x=3 my answer is 8 is that correct?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

thanx and you are quick

OpenStudy (anonymous):

via \[2^3=8\] so it is nice to be able to switch!

OpenStudy (anonymous):

you did the work. i just said "yes"

OpenStudy (anonymous):

one more \[\log_{5}125=x \]

OpenStudy (anonymous):

i have 25

OpenStudy (anonymous):

switch!

OpenStudy (anonymous):

no rewrite in exponential form

OpenStudy (anonymous):

write this in exponential form and the answer will be obvious. and obviously not 25 which is way way too big

OpenStudy (anonymous):

k give me a sec

OpenStudy (anonymous):

let me know what you get when you switch forms

OpenStudy (anonymous):

no cheating and looking at other post!

OpenStudy (anonymous):

3

OpenStudy (anonymous):

yes but i hope you realize why and why you need to be able to switch from one form to the other with ease

OpenStudy (anonymous):

takes two seconds to write \[5^x=125\] and then figure x = 3

OpenStudy (anonymous):

yes i am catching on but 9 years since any algebra doesnt make it very easy to come by....yes because 5^3 =125

OpenStudy (anonymous):

good

OpenStudy (anonymous):

not really much algebra to this yet though. just rewriting

OpenStudy (anonymous):

i guess